OpenStudy (lilai3):

Please help! Your help will be greatly appreciated cuz i have a test 2morro!!!! Will award with medal!!! We-R-Plastic makes plastic cylindrical garbage cans for kitchens and bathrooms. How much plastic is used for a can that has a diameter of 7 inches and is 14 inches tall, not including its lid? Round to the nearest tenth.

OpenStudy (anonymous):

Are you given a thickness or is it solid plastic?...

jimthompson5910 (jim_thompson5910):

why would you make a trash can that is completely solid and have no inner space to hold trash?

OpenStudy (lilai3):

i'm not really sure..that's why i need help...uhm...they didn't give a thickness. I typed every word and puncutation symbol from the problem my teacher gave me :P

jimthompson5910 (jim_thompson5910):

It sounds like all you need to do is find the volume of a cylinder with a diameter of 7 inches and height 14 inches

OpenStudy (lilai3):

@jim_thompson5910 i thought the same thing you did, but i thought this problem would be talking about surface arera, and not volume...unless i'm wrong..

jimthompson5910 (jim_thompson5910):

oh you're right, mixed up the two lol, my bad yeah you want to find out how much material you'll need to cover the outside so to speak

OpenStudy (lilai3):

so...it would be surface area right?

jimthompson5910 (jim_thompson5910):

so you would use SA = pi*r^2 + 2*pi*r*h keep in mind we're not adding the lid

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (comm.dan):

Yes surface area and there is a formula involved for surface area

OpenStudy (lilai3):

and the formula would be...?

jimthompson5910 (jim_thompson5910):

SA = pi*r^2 + 2*pi*r*h

OpenStudy (lilai3):

oh. thanks jim

OpenStudy (comm.dan):

A=BXH

OpenStudy (lilai3):

@Comm.Dan isn't it supposed to be a circle, so it's like sumthing to do with pi?

OpenStudy (comm.dan):

Yes it is @lilai3

OpenStudy (comm.dan):

I know it involves pi squared and 2 pi r

OpenStudy (anonymous):

Lateral area plus bottom lid:\[7 \pi 14+\pi \left(\frac{7}{2}\right)^2=346.4\text{ }\text{inch}^2 \]

OpenStudy (comm.dan):

Where in the world did you get that from @robtobey

OpenStudy (anonymous):

circumference of the cylinder times the cylinder height plus the area of the lower circle, "lower lid".

OpenStudy (comm.dan):

Alright Thx @robtobey

OpenStudy (anonymous):

You're welcome

OpenStudy (lilai3):

oh. thanks, all you guys, i wish i could give all of you a medal, and i'm sad that i can only choose one. thanks!

OpenStudy (anonymous):

@lilai3 Thank you for the medal.

OpenStudy (lilai3):

ur welcome. do you mind also helping me on my next question?